### v4

@@ 1 Introduction
The advent of the transformer architecture [1] and the introduction of the attention mechanism as its main computational component within the context of natural language processing (NLP) led to large advancements not only in language-related tasks but across all aspects related to the research and application of machine learning (ML). Transformers changed the landscape of NLP with the adoption of their architecture in designing highly successful and capable large language models (LLM) [2] such as GPT [3],LLama [4], Falcon [5] and Mistral [6]. The computer vision (CV) domain also experienced rapid adoption of transformer architectures. Vision-specific implementations such as ViT [7], MLP-Mixer [8], Conv-Mixer [9], and Swin Transformer [10] were introduced, along with many application-oriented designs that utilize such architectures, such as Detection Transformer (DETR) [11], Perceiver-IO [12], Unified-IO [13], DINO [14], and Segment Anything Model (SAM) [15].Given such wide adoption on many tasks and modalities, the need for more efficient implementations of the transformer has increased in importance, and in that direction, significant research attempts have been introduced, such as Linformer [16], FNets [17], Local-ViT [18], Max-ViT [19], and Nystromformer [20].The MLP-Mixer implementation of transformers is of interest as it introduces a unique design choice that targets a more efficient architecture with emphasis on the process of ”token mixing”. This mixing is applied in two stages [8]; the first stage applies the mixing within the input token representations, while the second stage applies the mixing process between corresponding positions in each token. A drawback of the MLP-Mixer design is that the mentioned mixing processes are performed with static weight matrices, which limits the capabilities of the architecture in comparison to the traditional transformers that utilize the dynamic process of the scaled dot product attention mechanism with sotmax activation function. At the same time, the traditional transformer architecture has its own drawback of quadratic complexity in input size [21], which imposes a considerable cost in both training and inference when selecting the architecture. It is notable that in the literature there is a lack of a design that adopts the efficiency measures introduced in the MLP-Mixer model while also maintaining a dynamic information filtering mechanism, as with the traditional transformer design. In this paper, we introduce a newly formulated computational block that can be used as a core process in constructing transformer architectures that blends both efficient elementary operations and dynamic information filtering. The new proposal utilizes the MLP-Mixer token mixing to learn a generator of dynamic per input gating function that selectively filters the input representation tokens that are then passed to the per token MLP stage as in traditional transformers, which results in a block that contains two levels of processing [22], an inner and an outer, hence the chosen name for the proposal as a Network in Network Transformer, or (NiNformer). In this work, the newly proposed architecture was trained and its performance evaluated with respect to multiple baselines that represent different architectural directions and a variety of design choices. The comparison was conducted on three datasets, and the experimentation was performed in an equalized setting with the same computational resources to ensure a fair evaluation. From the experiments conducted, it was observed that the NiNformer architecture was the most performing, and the obtained results verified the validity and capability of the underlying assumptions employed in our proposed computational block.
@@ 3.1 Baselines
Algorithm 1 overviews the ViT architecture.
Algorithm 1 ViT
@@
Algorithm 2 overviews the MLP-Mixer architecture.
Algorithm 2 MLP-Mixer
@@
Algorithm 3 overviews the Local-ViT architecture.
Algorithm 3 Local-ViT
@@ 3.2 Proposed Architecture
Algorithm 4 overviews our proposed NiNformer architecture.
Algorithm 4 NiNformer

### v5

@@ 1 Introduction
The advent of the transformer architecture [1] and the introduction of the attention mechanism as its main computational component within the context of natural language processing (NLP) led to large advancements not only in language-related tasks but across all aspects related to the research and application of machine learning (ML). Transformers changed the landscape of NLP with the adoption of their architecture in designing highly successful and capable large language models (LLM) [2] such as GPT [3],LLama [4], Falcon [5] and Mistral [6]. The computer vision (CV) domain also experienced rapid adoption of transformer architectures. Vision-specific implementations such as ViT [7], MLP-Mixer [8], Conv-Mixer [9], and Swin Transformer [10] were introduced, along with many application-oriented designs that utilize such architectures, such as Detection Transformer (DETR) [11], Perceiver-IO [12], Unified-IO [13], DINO [14], and Segment Anything Model (SAM) [15].Given such wide adoption on many tasks and modalities, the need for more efficient implementations of the transformer has increased in importance, and in that direction, significant research attempts have been introduced, such as Linformer [16], FNets [17], Local-ViT [18], Max-ViT [19], and Nystromformer [20].The MLP-Mixer implementation of transformers is of interest as it introduces a unique design choice that targets a more efficient architecture with emphasis on the process of ”token mixing”. This mixing is applied in two stages [8]; the first stage applies the mixing within the input token representations, while the second stage applies the mixing process between corresponding positions in each token. A drawback of the MLP-Mixer design is that the mentioned mixing processes are performed with static weight matrices, which limits the capabilities of the architecture in comparison to the traditional transformers that utilize the dynamic process of the scaled dot product attention mechanism with the softmax activation function. At the same time, the traditional transformer architecture has its own drawback of quadratic complexity in input size [21], which imposes a considerable cost in both training and inference when selecting the architecture. It is notable that in the literature there is a lack of a design that adopts the efficiency measures introduced in the MLP-Mixer model while also maintaining a dynamic information filtering mechanism, as with the traditional transformer design. In this paper, we introduce a newly formulated computational block that can be used as a core process in constructing transformer architectures that blends both efficient elementary operations and dynamic information filtering. The new proposal utilizes the MLP-Mixer token mixing to learn a generator of dynamic per input gating function that selectively filters the input representation tokens that are then passed to the per token MLP stage as in traditional transformers, which results in a block that contains two levels of processing [22], an inner and an outer, hence the chosen name for the proposal as a Network in Network Transformer, or (NiNformer). In this work, the newly proposed architecture was trained and its performance evaluated with respect to multiple baselines that represent different architectural directions and a variety of design choices. The comparison was conducted on three datasets, and the experimentation was performed in an equalized setting with the same computational resources to ensure a fair evaluation. From the experiments conducted, it was observed that the NiNformer architecture was the most performing, and the obtained results verified the validity and capability of the underlying assumptions employed in our proposed computational block.
@@ 3.1 Baselines
Procedure 1 overviews the ViT architecture.
Algorithm 1 Procedure 1 : ViT
@@
Procedure 2 overviews the MLP-Mixer architecture.
Algorithm 2 Procedure 2 : MLP-Mixer
@@
Procedure 3 overviews the Local-ViT architecture.
Algorithm 3 Procedure 3 : Local-ViT
@@ 3.2 Proposed Architecture
Procedure 4 overviews our proposed NiNformer architecture.
Algorithm 4 Procedure 4 : NiNformer